OpenStudy (anonymous):

Which of the following is not a step used to solve multi-step equations? A. identify values of variables in the equation from the written word problem B. finish by getting all variables and constants on the same side of the equation C. plug in known variables from the word problem into the equation D. translate the written phrases into an algebraic equation

OpenStudy (anonymous):

Write and simplify the equation you would use to solve the problem below. A farmer builds a rectangular pen with a perimeter of 116 feet. If the length is 5 feet more than the width, what is the width? A. w + 5 = 116 B. 4w + 10 = 116 C. 2w + 5 = 116 D. 2w + 10 = 116

OpenStudy (texaschic101):

B. finish by getting all variables and constants on the same side of the equation ======================= P = 2L + 2W L = W + 5 P = 116 116 = 2(W + 5) + 2W 116 = 2W + 10 + 2W 116 = 4W + 10 ===>answer B

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :)